OpenStudy (anonymous):

solve the DE \[{dy \over dx} = -{1 \over x + y}\]

OpenStudy (anonymous):

What is the latest method of differential equations are you learning?

OpenStudy (anonymous):

Try substituting in v = x + y -> dv/dx = dy/dx + 1.

OpenStudy (anonymous):

This becomes\[\frac{dv}{dx} - 1 = \frac{-1}{v}\]This is separable.

OpenStudy (anonymous):

The original question was find the member of the orthogonal trajectories of the family of curves x + y = ce^x - 1 that passes through the point (3, 0). I think that I did the first part right

OpenStudy (anonymous):

I got ln(v - 1) + v = x + c as an answer, with v = x + y.

OpenStudy (anonymous):

Ahh I don't know why this is so hard!! Well it's a first order differential equation so you might want to look over that section.

OpenStudy (anonymous):

\[\frac{dv}{dx} = \frac{v-1}{v} \rightarrow \int\limits \frac{v}{v-1}dv = \int\limits 1dx\]